Indicate the missing amount for each letter. Assume that in all cases manufacturing overhead is applied on the basis of direct abor cost and the rate is the same. (Round overhead rate to 2 decimal places, e.g. 15.25 and final answers to O decimal places, eg 5,275.) Case A Case B Case C $66,750 Direct materials used $ Direct labor Manufacturing $86,500 56,000 142,700 42,000 overhead applied Total manufacturing 149,750 213,750 costs Work in process 1/1/17 25,200 18,600 Total cost of work in procesS 207,500 G) Work in process 13,800 12/31/17 Cost of goods manufactured 199,300 228,000
Answer #1

Case ACase B $51,750$86,500$66,7510 Case C Direct Material used Direct Labor 56,0001,42,700 84,000 42,000 1,07,025 63 Manufacturing overhead applied Total Manufacturing costs 1,49,750 Work in process 1/1/17 Total cost of work in process Work in process 12/31/17 ,000 3,36,2252,13,750 25,200 18,600 2,07,5003,61,425 2,32,350 57,750 1:,a) 4,350 1,99,300 3,47,625 2,28,000 8,200 Cost of goods Manufactured

Over head rate is identified based on Case A Information ( i.e Direct labor and Manufacturing over head applied)

Over head rate = 42,000 / 56,000 * 100 = 75.00%

Case A :

a) Direct material used = Total manufacturing costs - (Direct Labor + Manufacturing over head applied)

Direct material used = 149,750 - (56,000 + 42,000)

Direct material used = $51,750

b) Work in process 1/1/17 = Total cost of work in process - Total manufacturing costs

Work in process 1/1/17 = 207,500 - 149,750

Work in process 1/1/17 = 57,750

c) Work in process 12/31/17 = Total cost of work in process - Cost of goods manufactured

Work in process 12/31/17 = 207,500 - 199,300

Work in process 12/31/17 = 8200

Case B :

d) Manufacturing over head applied = Direct labor * 75%

= 142,700 * 75/ 100

Manufacturing over head applied  = 107,025

e) Total manufacturing costs = Direct material used + Direct Labor + Manufacturing over head applied

= 86,500 + 142,700 + 107,025

Total manufacturing costs = 336,225

f) Total cost of work in process = Total manufacturing costs + Work in process 1/1/17

= 336,225 + 25,200

Total cost of work in process = 361,425

g) Cost of goods manufactured = Total cost of work in process -  Work in process 12/31/17

= 361,425 - 13,800

Cost of goods manufactured = 347,625